Three cards are chosen at random from a deck without replacement. What is the probability of getting a jack, a ten and a nine?
Answer by stanbon(75887)

Three cards are chosen at random from a deck without replacement. What is the probability of getting a jack, a ten and a nine?
---------------
P(jack) = 4/52
P(10) = 4/48
P(9) = 4/44
-------------------
P(j and t and n) = (4/52)(4/48)(4/44) = 64/109824 = 0.00058275...
Cheers,
Stan H.